The bank only stated the total number of certificates sold (125) and the total generated ($12,660). The bank manager needs to know the number of 6 month CODs purchased and the number of 12 month CODs purchased. 6 month CODs are $112 and 12 month CODs are $92. Write and solve a system of linear equations to determine how many of each type of certificate of deposits were sold.
Answer by Boreal(15235)

x=6 month
y=12 month
x+y=125
112x + 92y=12660
-92x-92y=-11500, multiplying the top by -92
112x+92=12660
20x=1160
x=58 6 month CD
y=67 12 month CD
58*112=$6496
67*92=$6194
sum is $12,660